# ACES: Automatic Cohort Extraction System for Event-Stream Datasets

### v1

@@ 2.4 Extraction Output
Figure 3: Sample output table from ACES. Continuing the example from Figure 2, we obtain this table after executing ACES extraction. The subject_id, index_timestamp, and label form the foundation of the defined cohort and are what the user would need, at a minimum, to obtain the necessary data for downstream tasks. In addition to these three columns, there exists a column for each of the defined windows in the task configuration, with its value being a structured object containing information about that particular window, such as the start and end times, as well as a summary of predicate counts. As seen in the values of this table, the timestamp of the end of the input window is always 24 hours after the trigger timestamp for each sample (row). Additionally, the window itself contains more than 5 of _ANY_EVENT—exactly matching our previously defined task configuration.
@@ 2.5 Python Usage
In addition to the command-line tool, we also provide a Python interface to allow researchers to easily leverage ACES for cohort extraction in their code pipelines. A full tutorial is provided at https://eventstreamaces.readthedocs.io/en/latest/notebooks/tutorial.html.
@@ 3 ACES Technical Design
    Structured Medical Data as Event-Stream Data: ACES views structured medical data as “event-stream” data, enabling the extraction of meaningful tasks and cohorts from continuous patient records.
@@
    Compatibility with Any Dataset: ACES is designed to be compatible with any dataset, provided the necessary dataset-specific predicates are extracted and formatted as an event-stream (this step is automated for any dataset that can be expressed in the MEDS [1] or ESGPT [10] formats) [^4] Note: This includes datasets that can be converted into these formats with existing ETLs, such as OMOP datasets through the MEDS OMOP ETL..
@@
The key principle behind the ACES algorithm is that we decompose the task of determining whether a section of a patient’s record meets certain criteria specified in a configuration file into a series of recursive analyses. These analyses evaluate whether specific windows, defined relative to the realizations of other windows in the patient record, meet their criteria. By imposing only minimal constraints on the structure of the task configuration file, we can ensure that tasks can always be extracted using this recursive approach, thus forming the foundation of the entire ACES system. For more details, please see the Appendix.
@@ 4 Using ACES: A Repository of Task Configuration Examples
To demonstrate the flexibility and utility of ACES, we define and publicly release the task configurations described in Table 1, both with dataset-agnostic criteria and with dataset-specific predicate realizations based on the MEDS version of the public MIMIC-IV dataset [4]. Additional usage profiles, tasks, and areas, as well as details of the task cohorts extracted from MIMIC-IV, can be found in the Appendix. These various tasks have been previously studied, and ACES will facilitate their conceptual reproducibility to encourage benchmarking efforts and robustness in ML for healthcare.
@@
readmission/30d Predict hospital readmission within 30 days of discharge.
mortality/post_hospital_discharge/30d Predict mortality within 30 days of discharge.
mortality/in_hospital/first_24h Predict mortality within a hospital admission using the first 24 hours of data from that admission.
mortality/in_hospital/first_48h Predict mortality within a hospital admission using the first 48 hours of data from that admission.
mortality/in_icu/first_24h Predict mortality within an ICU admission using the first 24 hours of data from that admission.
mortality/in_icu/first_48h Predict mortality within a ICU admission using the first 48 hours of data from that admission.
Table 1: A collection of sample configuration files for various common tasks on MIMIC-IV [4]. These tasks can be easily generalized to other datasets, such as e-ICU [15] or other private intensive care unit (ICU) and inpatient datasets. Please see the Appendix for further examples.
@@ 5 Discussion
Additional Related Work
@@
In addition to the existing tools discussed in Section 1, there are several other areas of related work relevant to ACES. Firstly, ACES could be directly connected with and support various health CDMs, such as OMOP, FHIR, PCORNET, and i2b2 [12, 2, 14, 11]. These models provide already-accepted standardized frameworks for organizing and analyzing healthcare data, and integrating ACES directly with them, rather than through ETLs, could greatly enhance its utility and interoperability. Static benchmarks that provide standardized datasets, evaluation metrics, and baseline methods for a range of clinical problems, such as YAIB [20], Harutyunyan et al.’s multitask learning clinical prediction benchmarks [3], and EHR-PT [8], can also be directly integrated with ACES to facilitate robust ML in healthcare. Lastly, ACES can be used in conjunction with various health data management tools, such as ESGPT [10], TemporAI [18], PyHealth [22], and OMOPLearn [7]. These tools offer functionalities for pre-processing, managing, and analyzing health data for downstream tasks, and integrating ACES with them directly can streamline ML workflows.
@@
Beyond healthcare, ACES is applicable to data from a variety of other domains, such as for financial, climate, or social media data—essentially, ACES could be used for any structured, longitudinal data that can be reformatted as an event-stream. This versatility makes ACES a powerful tool for extracting and analyzing complex event-based datasets across different fields.
Limitations & Future Roadmap
ACES has several key limitations that can be addressed in future work. Firstly, while already very expressive, ACES’s task configuration language is currently still relatively limited. Expressing more complex kinds of predicates, window aggregations, labeling functions, and criteria would expand the scope of ACES significantly. Similarly, the configuration language can also be made more user-friendly, such as by allowing users to specify a list of values that all establish a single predicate directly, rather than having to define multiple predicates for each value. ACES is also very well poised to capture more complex patterns of task and cohort relationships, including prescribed systems of case/control matching, automated analyses, or propensity re-weighting over excluded populations. It would also possible to enable users to nest ACES configuration files to leverage extracted task labels as new predicates in more complex ACES extraction processes. Finally, with the standardization that ACES offers, new opportunities for human-data interaction are also made available, such as via a natural language interface to define ACES predicates or configuration files and, thus, to extract downstream tasks, patient cohorts, or derived datasets in a code-free manner. Preliminary analyses of the viability of this process with current large language models show significant promise, though much future work remains.
ACES: Enabling a New Kind of Benchmark
@@
In addition to the clear impact of ACES on reproducibility, robustness, and accessibility of health datasets in ML for healthcare, we also feel that ACES is an essential tool for a “new kind of benchmark” in the field—and, in so being, is a portent of what needs to come should ML for healthcare progress to a more productive, communal, and impactful stage. In particular, we argue that for ML for healthcare to progress in the manner desired by the community, and most likely to be maximally positively impactful for all patients, we need to develop methodologies to test, share, and develop ML solutions across diverse datasets in a meaningful and reproducible manner, even without said datasets being publicly available to general researchers. This capability is critical because, without it, we will never be able to offer new inductive insights to clinical scientists or informaticians about which methods are most likely to work best on their novel, private data. In other words, if we cannot test our model training recipes across the diverse sets of clinical care settings, populations, and conceptual dataset schemas that exist in the real world, we similarly cannot expect those training recipes to generalize to said set of myriad downstream deployment areas. Tools like ACES, which make it as easy as possible for users to share the conceptual definitions of their tasks and prediction areas across datasets—in such a way that their colleagues can use them even over independent, private datasets—can help transform the kinds of benchmarking studies that we perform in ML for healthcare towards those that permit generalizable assessment of ML training recipes across datasets, clinical areas, and more.
@@ 6 Conclusion
In this work, we present the Automatic Cohort Extraction System for Event-Stream Datasets (ACES). ACES is a system designed to intuitively define cohorts and downstream tasks of interest in ML and reliably and automatically extract those cohorts from arbitrary datasets that are in an event-stream format. This system enables significantly greater shareability of task definitions, reproducibility of ML training and evaluation recipes, and is as easy to use as installing a package via pip and running a simple command-line tool. We feel that ACES will be an integral tool in the development of new kinds of benchmarks in ML for healthcare, which can be explored across both public and private datasets alike, and define and characterize the populations and tasks of interest in a manner that cleanly separates dataset-specific components from shareable dataset-agnostic components. To learn more about ACES and use it today in your work, please visit our GitHub repository at https://github.com/justin13601/aces and the online ACES documentation at https://eventstreamaces.readthedocs.io/en/latest/.
Acknowledgments and Disclosure of Funding
@@
Appendix
The full online ACES documentation is available at https://eventstreamaces.readthedocs.io/en/latest/. We have also included a compiled PDF version of this documentation as an ancillary file.
To answer specific questions about ACES, please see the below index (links to the external online documentation are provided; references to chapters and page numbers of the ancillary PDF documentation are in brackets):
@@
    Sample task descriptions and specifications are provided on the Task Examples page of the online documentation. Please refer to https://eventstreamaces.readthedocs.io/en/latest/notebooks/examples.html (Chapter 3, Pg 21-26).
@@
      For an overview of predicates and how they form the foundation of ACES, please refer to the Predicates DataFrame page at https://eventstreamaces.readthedocs.io/en/latest/notebooks/predicates.html (Chapter 4, Pg 27-30).
@@
      A window in ACES is defined at https://eventstreamaces.readthedocs.io/en/latest/algorithm.html#window. Additionally, for details on how to define a window, please refer to https://eventstreamaces.readthedocs.io/en/latest/readme.html#windows (Chapter 1.4.3, Pg 9-10).
@@
    For general ACES usage instructions, please refer to https://eventstreamaces.readthedocs.io/en/latest/usage.html#quick-start (Chapter 2.1, Pg 13-16). Additionally, brief end-to-end instructions are also available at https://eventstreamaces.readthedocs.io/en/latest/readme.html#instructions-for-use (Chapter 1.3, Pg 4-6).
@@
      For detailed instructions on using ACES CLI, please refer to the online Usage Guide at https://eventstreamaces.readthedocs.io/en/latest/usage.html#detailed-instructions (Chapter 2.2, Pg 16-17).
@@
      For a step-by-step tutorial on using the ACES Python API, please refer to the online Code Example Notebook at https://eventstreamaces.readthedocs.io/en/latest/notebooks/tutorial.html (Chapter 5, Pg 31-41).
@@
    ACES with and vs. Other Tools
    For an overview of how ACES could be used with other existing complementary tools for reproducible machine learning, please refer to the online section at https://eventstreamaces.readthedocs.io/en/latest/readme.html#complementary-tools (Chapter 1.5.2, Pg 10).
    For an overview of ACES and other existing alternative tools for semi- or fully-automated cohort extraction, please refer to the online section at https://eventstreamaces.readthedocs.io/en/latest/readme.html#alternative-tools (Chapter 1.5.3, Pg 10).
@@
    For technical details on the ACES configuration language, please refer to the Configuration Language Specification page at https://eventstreamaces.readthedocs.io/en/latest/configuration.html (Chapter 6.1, Pg 43-46).
@@
    For a glossary of terminology used throughout ACES, please refer to the Algorithm Terminology page at https://eventstreamaces.readthedocs.io/en/latest/technical.html#algorithm-terminology (Chapter 6.3, Pg 49-50).
@@
    For technical details on the ACES algorithm, please refer to the Algorithm Design page at https://eventstreamaces.readthedocs.io/en/latest/technical.html#algorithm-design (Chapter 6.4, Pg 50-53).
@@
    For the complete ACES module documentation, including tests that ensure algorithm correctness, please refer to the Module API page at https://eventstreamaces.readthedocs.io/en/latest/api/modules.html (Chapter 8, Pg 57-114).
@@
    For an overview of the computational profile of ACES, please refer to the Computational Profile page at https://eventstreamaces.readthedocs.io/en/latest/profiling.html (Chapter 7, Pg 55-57).
@@
    For additional examples of configuration files and criteria of different machine learning for healthcare tasks, please refer to https://github.com/mmcdermott/PIE_MD/tree/main/tasks/criteria.

### v2

@@ 2.5 Python Usage
In addition to the command-line tool, we also provide a Python interface to allow researchers to easily leverage ACES for cohort extraction in their code pipelines. A full tutorial is provided at: https://eventstreamaces.readthedocs.io/en/latest/notebooks/tutorial_meds.html. Please also refer to the PDF compilation of the tutorial notebook in the Supplementary Material (Pages 31-36).
@@ 3 ACES Technical Design
    Structured Medical Data as Event-Stream Data: ACES views structured medical data as “event-stream” data1, enabling the extraction of meaningful tasks and cohorts from continuous patient records.
@@
    Compatibility with Any Dataset: ACES is designed to be compatible with any dataset, provided the necessary dataset-specific predicates are extracted and formatted as an event-stream (this step is automated for any dataset that can be expressed in the MEDS (Arnrich et al., 2024) or ESGPT (McDermott et al., 2023) formats) [^4] Note: This includes datasets that can be converted into these formats using existing ETLs, such as OMOP datasets through the MEDS OMOP ETL..
@@
The key novel principle behind the ACES algorithm is that we decompose the task of determining whether a section of a patient’s record meets certain criteria specified in a configuration file into a series of recursive analyses. These analyses evaluate whether specific windows, defined relative to the realizations of other windows in the patient record, meet their criteria. By imposing only minimal constraints on the structure of the task configuration file, we can ensure that tasks can always be extracted using this recursive approach, thus forming the foundation of the entire ACES system. For more details, please see the Appendix and the technical documentation in the Supplementary Materials.
Figure 3: Sample output table from ACES. Continuing the example from Figure 2, we obtain this table after executing ACES extraction. The subject_id, index_timestamp, and label form the foundation of the defined cohort and are what the user would need, at a minimum, to perform downstream tasks. In addition to these three columns, there exists a column for each of the defined windows in the task configuration, with its value being a structured object containing information about that particular window, such as the start and end times, as well as a summary of predicate counts. As seen in the values of this table, the timestamp of the end of the input window is always 24 hours after the trigger timestamp for each sample (row), and contains more than 5 of _ANY_EVENT—exactly matching our previously defined task configuration.
@@ 4 Using ACES: A Repository of Task Configuration Examples
To demonstrate the flexibility and utility of ACES, we define and publicly release the task configurations described in Table 1, both with dataset-agnostic criteria and with dataset-specific predicate realizations based on the MEDS version of the public MIMIC-IV dataset. Additional usage profiles, tasks, and areas, as well as details of the task cohorts extracted from MIMIC-IV, can be found in the Appendix and documentation. These various tasks have been previously studied, and ACES will facilitate their conceptual reproducibility to encourage benchmarking efforts and ensure robustness in ML for healthcare.
Table 1: A collection of sample configuration files for various common predictive tasks on MIMIC-IV (Johnson et al., 2023). These tasks can be easily generalized to other datasets, such as e-ICU (Pollard et al., 2018) or other private intensive care unit (ICU) and inpatient datasets by simply swapping out appropriate predicate definitions. pastecode.io links have been provided.
@@
First 24h in-hospital mortality Predict mortality within a hospital admission using the first 24 hours of data from that admission.
First 48h in-hospital mortality Predict mortality within a hospital admission using the first 48 hours of data from that admission.
First 24h in-ICU mortality Predict mortality within an ICU admission using the first 24 hours of data from that admission.
First 48h in-ICU mortality Predict mortality within a ICU admission using the first 48 hours of data from that admission.
30d post-hospital-discharge mortality Predict mortality within 30 days of discharge.
30d re-admission Predict hospital readmission within 30 days of discharge.
Myocardial infarction 1-5Y phenotyping Predict myocardial infarction (MI) incidence 1-5 years after hospital admission.
Reduced echo-derived LVEF 9m post-ECG Predict reduced echo-derived left ventricular ejection fraction (LVEF) within 9 months of any ECG.
CKD onset in diabetics 5Y from kidney panel Predict chronic kidney disease (CKD) onset in diabetic patients within 5 years of any kidney panel laboratory test.
4.1 Computational Profile
To establish an overview of the computational profile of ACES, a collection of various common tasks was queried on MIMIC-IV (Johnson et al., 2023) in the MEDS format. The MIMIC-IV MEDS schema has approximately 50,000 patients per shard with an average of approximately 80,500,000 total event rows per shard over seven shards.
Table 2: Performance statistics for various common predictive tasks on a single MEDS shard of MIMIC-IV (Johnson et al., 2023), including the number of patients, the number of samples, total runtime (in seconds), and maximum memory usage (in MiBs).
Task # Patients # Samples Total Time (s) Max Memory (MiBs)
First 24h in-hospital mortality 20,971 58,823 363.09 106,367.14
First 48h in-hospital mortality 18,847 60,471 364.62 108,913.95
First 24h in-ICU mortality 4,768 7,156 216.81 39,594.37
First 48h in-ICU mortality 4,093 7,112 217.98 39,451.86
30d post-hospital-discharge mortality 28,416 68,547 182.91 30,434.68
30d re-admission 18,908 464,821 367.41 106,064.04
Myocardial Infarction 1-5Y phenotyping 3,329 8,319 198.04 33,427.70
Reduced echo-derived LVEF 9m post-ECG 14 17 210.02 35,385.79
CKD onset in diabetics 5Y from kidney panel 736 3,503 238.65 44,221.81
All experiments were executed on a Linux server with 36 cores and 340 GBs of RAM available. A single MEDS shard was used to provide a bounded computational overview of ACES, which is applicable even when scaled to larger datasets using Hydra3. For instance, if one shard costs $M$ memory and $T$ time, then $N$ shards may be executed in parallel with $N*M$ memory and $T$ time, or in series with $M$ memory and $T*N$ time.
@@ 5 Discussion
5.1 Additional Related Work
@@
In addition to the existing tools discussed in Section 1, there are several other areas of related work relevant to ACES. Firstly, ACES could be directly connected with and support various health CDMs, such as OMOP, FHIR, PCORNET, and i2b2 (OHDSI, 2021; Bender & Sartipi, 2013; PCORnet, 2023; Murphy et al., 2010), rather than through ETLs. These models provide already-accepted standardized frameworks for organizing and analyzing healthcare data, and integrating ACES directly with them could greatly enhance its utility and interoperability. Similarly, frameworks such as DescEmb (Hur et al., 2022) and GenHPF (Hur et al., 2024) hold great synergistic potential with ACES, and we believe that they can be complementary in enabling new kinds of cross-dataset training, transfer learning, and evaluation. Static benchmarks that provide standardized datasets, evaluation metrics, and baseline methods for a range of clinical problems, such as YAIB (van de Water et al., 2024), Harutyunyan et al.’s multitask learning clinical prediction benchmarks (Harutyunyan et al., 2019), and EHR-PT (McDermott et al., 2021a), can also be directly integrated with ACES to facilitate robust ML in healthcare. Lastly, ACES can be used in conjunction with various health data management tools, such as ESGPT (McDermott et al., 2023), TemporAI (Saveliev & van der Schaar, 2023), PyHealth (Yang et al., 2023), and OMOPLearn (Kodialam et al., 2021). These tools offer functionalities for pre-processing, managing, and analyzing health data for downstream tasks, and integrating ACES with them directly can streamline ML workflows.
@@
Beyond healthcare, ACES is applicable to data from a variety of other domains, such as for financial, climate, or social media data—essentially, ACES could be used for any structured, longitudinal data that can be reformatted as an event-stream. This versatility makes ACES a powerful library for extracting and analyzing complex event-based datasets across different fields.
5.2 Limitations & Future Roadmap
ACES has several key limitations that can be addressed in future work. Firstly, while already very expressive, ACES’s task configuration language can still be further expanded. Expressing more complex kinds of predicates, window aggregations, labeling functions, and criteria would expand the scope of ACES significantly. ACES also seeks to provide direct support for cohort extraction based on unstructured data (notes and memos) in the future, though currently, such predicates need to be manually extracted by the user.
ACES is also very well poised to capture more complex patterns of task and cohort relationships, including prescribed systems of case/control matching, automated bias analyses, or propensity re-weighting over excluded populations. It is also possible to enable users to nest ACES configuration files to leverage extracted task labels as new predicates in more complex ACES extraction processes.
To enhance the scalability of ACES, we will ensure that the expanding MEDS standard will always be supported. Direct interoperability with existing resources in this space, in particular ATLAS and OHDSI vocabulary-derived cohort definitions, is a high priority area for future work.
Finally, with the standardization that ACES offers, new opportunities for human interaction with data are also made available, such as via a natural language interface to define ACES predicates or configuration files and, thus, to extract downstream tasks, patient cohorts, or derived datasets in a code-free manner on diverse input EHR formats. We aim to also explore the viability of leveraging large language models (LLMs) to directly format predicate and criteria definitions given a data dictionary; preliminary work on this front shows very promising results!
5.3 ACES for Enabling a New Kind of Benchmark
In addition to the clear impact of ACES on reproducibility, robustness, and accessibility of health datasets and ML for healthcare, we also feel that ACES is critical for a “new kind of benchmark” in the field—and, in so being, is a portent of what needs to come should ML for healthcare progress to a more productive, communal, and impactful stage.
@@ 5.3 ACES for Enabling a New Kind of Benchmark
In particular, we argue that for this field to progress in the manner desired by the community, and most likely to be maximally positively impactful for all patients, we need to develop methodologies to test, share, and develop ML solutions across diverse datasets in a meaningful and reproducible manner, even without said datasets being publicly available to general researchers. This capability is critical because, without it, we will never be able to offer new inductive insights about which methods are most likely to work best on novel, private data. In other words, if we cannot test our model training recipes across the diverse sets of clinical care settings, populations, and conceptual dataset schemas that exist in the real world, we similarly cannot expect those training recipes to generalize to said set of myriad downstream deployment areas.
Libraries like ACES, which make it as easy as possible for users to share the conceptual definitions of their tasks and prediction areas across datasets—in such a way that their colleagues can use them even over independent, private datasets—can help transform the kinds of benchmarking studies we can perform in ML for healthcare towards those that permit generalizable assessment of ML training recipes across datasets, clinical areas, and more.
@@ 6 Conclusion
In this work, we present the Automatic Cohort Extraction System for Event-Stream Datasets (ACES). ACES is a system designed to intuitively define cohorts and downstream tasks of interest for learning representations and reliably and automatically extract those cohorts from arbitrary datasets that are in an event-stream format. This system enables significantly greater shareability of task definitions, reproducibility of ML training and evaluation recipes, and is as easy to use as installing a package via pip and running a simple command-line tool. We feel that ACES will be integral in the development of new kinds of benchmarks in ML for healthcare, which can be explored across both public and private datasets alike, and characterize the populations and tasks of interest in a manner that cleanly separates dataset-specific components from shareable dataset-agnostic components. To learn more about ACES and use it today in your work, please visit our GitHub repository at: https://github.com/justin13601/aces, and the online ACES Documentation at: https://eventstreamaces.readthedocs.io/en/latest/index.html.
7 Video Demonstration
To complement this paper and the documentation, we have provided a very short online video demonstration of ACES in action, available here: https://youtu.be/i_hCaHDydqA. It overviews the Python API and walks through a running example of extracting a cohort for an in-hospital mortality prediction task from a synthetic MEDS dataset. Additionally, it briefly covers several simple and complex real-world task configuration files.
8 Reproducibility Statement
We have provided links to the ACES library source code and pastecode.io snippets of configuration files used in the experiments. Reproducible tutorial notebooks and technical implementation details are included in the Supplementary Materials (Pages 31-36 and 41-51, respectively) and are accessible via the online public documentation at: https://eventstreamaces.readthedocs.io/en/latest/.
9 Acknowledgments and Disclosure of Funding
@@
Appendix A Documentation
The full ACES Online Documentation is available at: https://eventstreamaces.readthedocs.io/en/latest/. We have also included a compiled PDF version of this documentation in the Supplementary Material.
To answer specific questions about ACES, please see the below index (links to the external online documentation, as well as references to chapters and page numbers of the ancillary PDF documentation are provided):
@@ Appendix A Documentation
    Sample task descriptions and specifications are provided on the Task Examples page of the online documentation (https://eventstreamaces.readthedocs.io/en/latest/notebooks/examples.html). Please refer to Chapter 3, Pages 21-26.
@@
      For an overview of predicates and how they form the foundation of ACES, please refer to the Predicates DataFrame section (https://eventstreamaces.readthedocs.io/en/latest/notebooks/predicates.html) in Chapter 4, Pages 27-30.
@@
      A window in ACES is defined as a segment in the patient record (https://eventstreamaces.readthedocs.io/en/latest/algorithm.html#window). For details on how to define a window (https://eventstreamaces.readthedocs.io/en/latest/readme.html#windows), please refer to Chapter 1.3.3, Page 9.
@@
    For general ACES usage instructions (https://eventstreamaces.readthedocs.io/en/latest/usage.html#quick-start), please refer to Chapter 2.1, Pages 13-16. Additionally, brief end-to-end instructions (https://eventstreamaces.readthedocs.io/en/latest/readme.html#instructions-for-use) are also available in Chapters 1.2 and 1.3, Pages 4-10.
@@
      For detailed instructions on using ACES CLI (https://eventstreamaces.readthedocs.io/en/latest/usage.html#detailed-instructions), please refer to the Usage Guide in Chapter 2.2, Pages 16-20.
@@
      For a step-by-step tutorial on using the ACES Python API (https://eventstreamaces.readthedocs.io/en/latest/notebooks/tutorial_meds.html), please refer to the Code Example Notebook in Chapter 5, Pages 31-41.
@@
    ACES with & vs. Other Tools
    For an overview of how ACES could be used with other existing complementary tools for reproducible machine learning (https://eventstreamaces.readthedocs.io/en/latest/readme.html#complementary-tools), please refer to Chapter 1.4.2, Page 10.
    For an overview of ACES and other existing alternative tools for semi- or fully-automated cohort extraction (https://eventstreamaces.readthedocs.io/en/latest/readme.html#alternative-tools), please refer to Chapter 1.4.3, Page 11.
@@
    For technical details on the ACES configuration language (https://eventstreamaces.readthedocs.io/en/latest/configuration.html), please refer to the Configuration Language Specification section in Chapter 6.1, Pages 41-44.
@@
    For a glossary of terminology used throughout ACES (https://eventstreamaces.readthedocs.io/en/latest/technical.html#algorithm-terminology), please refer to the Algorithm Terminology section in Chapter 6.3, Pages 46-47.
@@
    For technical details on the ACES algorithm (https://eventstreamaces.readthedocs.io/en/latest/technical.html#algorithm-design), please refer to the Algorithm Design section in Chapter 6.4, Pages 48-51.
@@
    For the complete ACES module documentation (https://eventstreamaces.readthedocs.io/en/latest/api/modules.html), including doctests that ensure algorithm correctness, please refer to the Module API sections in Chapter 8, Pages 55-125.
@@
    For an overview of the computational profile of ACES (https://eventstreamaces.readthedocs.io/en/latest/profiling.html), please refer to the Computational Profile section in Chapter 7, Page 53.
@@
    For additional examples of configuration files and criteria of different machine learning for health tasks, please refer to the MEDS-DEV benchmarking effort on GitHub (https://github.com/mmcdermott/MEDS-DEV/tree/main).